OpenStudy (anonymous):

When I get P(Z < -1.61), how should I proceed to find the probability using table P(Z

OpenStudy (anonymous):

I was thinking P (Z < -1.61) = 1 - P( Z < 1.61) = xxx Does that make sense?

OpenStudy (dumbcow):

yes that will work

OpenStudy (dumbcow):

depends on table you use if table has neg Z values, then just look up Z=-1.61

OpenStudy (anonymous):

What if P(Z < - 51)?

OpenStudy (dumbcow):

that is essentially 0....about 99% of distribution is within 3 std dev

OpenStudy (anonymous):

I also got 0, but couldn't believe it :( Thanks!!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

Just to make sure I was working correctly: \[\mu = 40\]\[\sigma^2=38.4\]\[\sqrt{n}=\sqrt{1000}\] Find P(X<30) \[Z = \frac{30-40}{\sigma/\sqrt{n}}=-51\]P(X<30) ~= P (Z<-51) = 1-P(51.03) = 1-1 =0?

OpenStudy (anonymous):

@dumbcow Sorry for interrupting :(

OpenStudy (dumbcow):

no prob didn't see you posted

OpenStudy (anonymous):

Could you please check to see if my work makes sense?

OpenStudy (dumbcow):

is the variance a "sample variance" ?

OpenStudy (anonymous):

I found the variance by \[\sigma^2 = npq\]

OpenStudy (dumbcow):

ok seems like it all checks out :) due to large sample size, the Z value is very high

OpenStudy (anonymous):

Is 1000 a large sample size?

OpenStudy (dumbcow):

yes , usually anything over 50 is considered large or "significant"

OpenStudy (anonymous):

Wow! Thanks a ton!!! :)